# PanopticNeRF-360: Panoramic 3D-to-2D Label Transfer in Urban Scenes

### v3

@@ Abstract
3D-to-2D Label Transfer, Panoptic Labeling, Neural Rendering, Urban Scene Understanding
@@ 3.1 Problem Formulation
Let $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ denote a camera ray. The color at the corresponding pixel can be obtained by volume rendering [52]
$\mathbf{C}(\mathbf{r})=\sum_{i=1}^{N}T_{i}(1-\exp(-\sigma_{i}\delta_{i}))\mathbf{c}_{i}\ ,\ T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)$ (3)
where $\sigma_{i}$ and $\mathbf{c}_{i}$ are density and color value at point $i$ sampled along the ray, $T_{i}$ denotes the transmittance at the sample point, and $\delta_{k}=t_{i+1}-t_{i}$ is the distance between adjacent samples. Let $\pi$ denote the volume rendering process of a ray $\mathbf{r}$, Eq. 3 can be rewritten as $\mathbf{C}(\mathbf{r}|\boldsymbol{\theta})=\pi(\mathbf{c})$.
@@ 3.4 Joint Geometry and Semantic Optimization
3.5 Implementation Details
Training: We optimize one PanopticNeRF-360 model per scene, using a single NVIDIA RTX 3090. For each scene, we set the origin to the center of the scene. We use Adam [35] with a learning rate of 5e-4 to train our models. We set the latent appearance code length to $n=12$, and loss weights to $\lambda_{\hat{\mathbf{S}}}=2,\lambda_{{\mathbf{S}}}=1,\lambda_{{\mathbf{s}}}=1,\lambda_{\hat{\mathbf{T}}}=2,\lambda_{\mathbf{C}}=1,\lambda_{d}=0.1$, and the density threshold to $\sigma_{th}=0.1$. We optimize the total loss $\mathcal{L}$ for 30,000 iterations. For the second stage of refining instances, we further fine-tune the model for 4,000 iterations.
4 Experiments
Metrics: Following  [19], we evaluate semantic labels via the mean Intersection over Union (mIoU) and the average pixel accuracy (Acc) metrics. To quantitatively evaluate multi-view consistency (MC), we utilize LiDAR points to retrieve corresponding pixel pairs between two consecutive evaluation frames. The MC metric is then calculated as the ratio of pixel pairs with consistent semantic labels over all pairs. For evaluating panoptic segmentation, we report Panoptic Quality (PQ) [36], which can be decomposed into Segmentation Quality (SQ) and Recognition Quality (RQ). We additionally adopt PQ† [62] as PQ over-penalizes errors of stuff classes. To evaluate perspective and fisheye labels in a unified manner, we design ${metric}^{\ast}$ (e.g., mIoU* and PQ*) that reports the average of perspective metric and fisheye metric. To verify that PanopticNeRF-360 is able to improve the underlying geometry, we further evaluate on perspective views the rendered depth compared to sparse depth maps obtained from LiDAR using Root Mean Squared Error (RMSE) and the ratio of accurate predictions ($\delta_{1.25}$[3, 18].
@@ 4 Experiments
4.1 Label Transfer
As most baselines are not designed for panoptic label transfer, we first compare the semantic predictions of all methods and then compare the panoptic predictions to 3D-2D CRF.
Semantic Label Transfer: As shown in Table I, Fig. 8 and Fig. 9, our method achieves the highest mIoU and Acc quantitatively and qualitatively. Specifically, compared to 3D-2D CRF, we obtain an absolute mIoU improvement of $2.4\%$ on perspective views and $11.9\%$ on fisheye views. The performance gap widens on the fisheye view as the pseudo-GT is less accurate on fisheye imagery. This observation also highlights the value of our method which is able to provide RGB images and semantic labels at novel viewpoints to enhance the generalization ability of 2D perception models. Despite PSPNet* being finetuned on KITTI-360 which reduces the performance gap, our method outperforms PSPNet* by a large margin. Supervised by the extremely sparse manually annotated GT, Semantic NeRF struggles to produce reliable performance. Using pseudo labels of PSPNet* and Tao et al., JacobiNeRF and Panoptic Lifting are capable of denoising and thus improving performance $(67.2\%\rightarrow 68.5\%/69.6\%,46.3\%\rightarrow 52.1\%/52.5\%)$. However, both JacobiNeRF and Panoptic Lifting are inferior in urban scenarios when the input views are sparse. In terms of perspective MC, ours is comparable to 3D-2D CRF and significantly surpasses 2D-to-2D label transfer methods. While our method slightly lags behind 3D Point + GC in terms of MC, this can be explained by the fact that label consistency is evaluated on sparsely projected 3D points which GC takes as input to generate a dense label map.
Panoptic Label Transfer: We compare against Panoptic Lifting and 3D-2D CRF for panoptic label transfer, as the other baselines do not support distinguishing instances. As Panoptic Lifting is limited to segment instance classes known to pre-trained 2D panoptic segmentation models, it is not capable of segmenting the “building” class. Therefore, we consider two settings for quantitatively measuring the panoptic segmentation metrics: “building” as stuff when comparing with Panoptic Lifting, and “building” as thing when comparing with 3D-2D CRF. Table II shows that PanopticNeRF-360 outperforms the baselines in both cases. As shown in Fig. 10, Panoptic Lifting can not reconstruct instances well in far regions where Mask2Former and their instance assignment strategy perform poorly, yielding a larger gap in terms of the panoptic segmentation metrics. In contrast, PanopticNeRF-360 avoids this issue by leveraging weak 3D labels, consistently producing good performance in both near and far regions. Our proposed method also outperforms the 3D-2D CRF in both things and stuff classes, depsite that 3D-2D CRF leverages the same weak 3D labels as input. Visual comparisons are shown in Fig. 10 and Fig. 11. As can be seen, our method gracefully handles over- and under-exposure at buildings, which is a challenge for the 3D-2D CRF, as it projects LiDAR points to reconstruct the intermediate meshes whose quality suffers in these scenarios (see Supplementary).
@@ 4.1 Label Transfer
4.2 Neural Scene Representation
@@
TABLE III: Neural Scene Representation Study over 4 scenes. We bold and underline the two best performing methods, respectively.
@@
We further compare iNGP [54], Tri-planes [9] (TensoRF-VM), and MLP [52] as scene representation. Towards this goal, we analyze the results over 4 scenes where ambient conditions vary between each other. For a fair comparison, we run each scene for 30,000 iterations. Please refer to the supplementary for implementation details. Table III indicates that while the pure MLP demonstrates superior label quality, it falls short in appearance reconstruction. On the other hand, iNGP excels in novel view synthesis ($\sim$3.5dB higher than MLP) and convergence, albeit at the cost of subpar semantic label quality in comparison to the MLP. Our proposed scene feature, a hybrid of MLP and hash grids, achieves a tradeoff between label quality and appearance. Tri-planes fail to represent high-fidelity scene structure both in terms of semantic labels and appearance, especially at boundaries between foreground and background (sky) and in over-exposed regions as grid-based methods are not robust to filter noise and tend to suffer from local minimum. While iNGP is also grid-based, it conditions features on multi-scale levels, thereby potentially aggregating global information, and employs a hash function for non-periodic signal embedding. As depicted in Fig. 13, iNGP contains more noise than ours, and Tri-planes fail to reconstruct the overall label, especially at buildings over-exposed by sunlight.
@@ 4.3 Synthesized Labels for Perception Models
To validate that the synthesized novel view labels are important for autonomous driving and robotics applications, we fine-tune Mask2Former [13] pre-trained on CityScape [15] using our synthesized semantic labels on the KITTI-360. We fine-tune it using our synthesized semantic labels on the KITTI-360. Specifically, we divide 10 test scenes into 8 for training and 2 for test. For each scene, we render 3 semantic maps at each of 64 trajectory points using the left perspective camera: one in the original direction, one rotated horizontally 30° to the left, and another rotated horizontally 30° to the right, resulting in a total of 1,536 training samples. During fine-tuning, we format the data to match CityScape’s and exclude categories not present in CityScape. We fine-tune the model for 5 epochs with a batch size of 8, using a learning rate of 0.0001 and AdamW optimizer. As illustrated in Table IV and Fig. 15, the fine-tuned Mask2Former (Mask2Former-ft) significantly outperforms the original (+3.7 mIoU on perspective views, and +6.3 mIoU on fisheye views), especially on Sidewalk and Wall. This improvement demonstrates the effectiveness of our synthesized labels in enhancing the perception model’s performance. Moreover, the larger improvement on the fisheye views further verifies that existing perception models degenerates on unseen viewpoints and demonstrates the importance of being able to synthesizing labels with large viewpoint changes.
4.4 Label Transfer on Waymo Dataset
We additionally conduct experiments on the Waymo dataset [50] to showcase PanopticNeRF-360’s generalization ability. We choose two scenes, each consisting of 198 consecutive frames with the front-view camera. During data preparation, we use the KITTI-360 annotation toolkit to label 3D bounding primitives and generate 2D pseudo semantic labels with Mask2Former [13]. This annotation takes around 2 hours for each scene, i.e., the average per-frame labeling time takes only 0.6 minutes. We illustrate the entire set of labeled bounding boxes and camera trajectory in Fig. 14. Ours not only renders high-fidelity appearance (28.28 dB on test views), but also generates high-quality semantic and instance masks on both pre-recorded and novel viewpoints. This demonstrates the applicability of our label transfer method.
Figure 14: Qualitative Label Transfer Results on Waymo. Our method can render high-quality panoptic label maps as well as high-frequence appearance on the Waymo dataset.
Method Test View Road Sdwlk Terr Bldg Vegt Car Wall Fence Sky mIoU Acc
Mask2Former Perspective 93.1 61.3 37.0 88.7 87.7 92.3 31.8 61.4 94.8 72.0 89.3
Mask2Former-ft 94.6 72.1 38.4 90.9 89.3 93.4 40.9 66.8 94.6 75.7 91.1
Mask2Former Fisheye 84.5 60.7 20.5 80.9 77.1 60.9 26.5 42.3 98.4 61.3 89.1
Mask2Former-ft 88.4 70.3 27.2 84.9 88.4 62.3 43.1 45.8 98.3 67.6 90.8
TABLE IV: Quantitative Comparison of Semantic Label on 2 test scenes of KITTI-360.
@@ 4.6 Limitations
5 Conclusion
We present PanopticNeRF-360 that infers in 3D space and renders omnidirectional per-pixel semantic and instance labels for 3D-to-2D label transfer. By unifying coarse 3D bounding primitives and noisy 2D semantic predictions, PanopticNeRF-360 is capable of performing mutual enhancement of geometry and semantics compared to naïve joint optimization of geometry and semantics. Specifically, it improves the underlying scene geometry given sparse input views leveraging label-guided geometry optimization, while concurrently resolving label noise based on improved geometry. Moreover, it enables label synthesis at a large range of novel viewpoints, including panoramic perspectives. We posit that our method marks a significant step towards improving data annotation efficiency while delivering a consistent, continuous 3D panoptic representation of a scene.

### v4

@@ Abstract
3D-to-2D Label Transfer, Panoptic Labeling, Semantic Labeling, Neural Rendering, Urban Scene Understanding
@@ 3.1 Problem Formulation
[…]
$\noindent{\color[rgb]{0,0,0}{\mathbf{C}(\mathbf{r})=\sum_{i=1}^{N}\mathcal{T}_{i}(1-\exp(-\sigma_{i}\delta_{i}))\mathbf{c}_{i}\ ,\ \mathcal{T}_{i}=\exp}}\noindent{\color[rgb]{0,0,0}{\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)}}$ (3)
where $\sigma_{i}$ and $\mathbf{c}_{i}$ are density and color value at point $i$ sampled along the ray, $\noindent{\color[rgb]{0,0,0}{\mathcal{T}_{i}}}$ denotes the transmittance at the sample point, and $\delta_{j}=k_{i+1}-k_{i}$ is the distance between adjacent samples. Let $\pi$ denote the volume rendering process of a ray $\mathbf{r}$, Eq. 3 can be rewritten as $\mathbf{C}(\mathbf{r}|\boldsymbol{\theta})=\pi(\mathbf{c})$.
@@ 3.4 Joint Geometry and Semantic Optimization
3.5 Implementation Details
Training: We optimize one PanopticNeRF-360 model per scene, using a single NVIDIA RTX 3090. For each scene, we set the origin to the center of the scene. We use Adam [35] with a learning rate of 5e-4 to train our models. We set the latent appearance code length to $n=12$, and loss weights to $\lambda_{\hat{\mathbf{S}}}=2,\lambda_{{\mathbf{S}}}=1,\lambda_{{\mathbf{s}}}=1,\lambda_{\hat{\mathbf{T}}}=2,\lambda_{\mathbf{C}}=1,\lambda_{d}=0.1$, and the density threshold to $\sigma_{th}=0.1$. We optimize the total loss $\mathcal{L}$ for 30,000 iterations. For the stage of refining instances, we further fine-tune the model for 4,000 iterations.
4 Experiments
@@
Figure 13: Qualitative Label Transfer Results on Waymo. Our method can render high-quality panoptic label maps as well as high-frequency appearance on the Waymo dataset.
Metrics: Following  [19], we evaluate semantic labels via the mean Intersection over Union (mIoU) and the average pixel accuracy (Acc) metrics. To quantitatively evaluate multi-view consistency (MC), we utilize LiDAR points to retrieve corresponding pixel pairs between two consecutive evaluation frames. The MC metric is then calculated as the ratio of pixel pairs with consistent semantic labels over all pairs. For evaluating panoptic segmentation, we report Panoptic Quality (PQ) [36], which can be decomposed into Segmentation Quality (SQ) and Recognition Quality (RQ). We additionally adopt PQ† [62] as PQ over-penalizes errors of stuff classes. To evaluate perspective and fisheye labels in a unified manner, we design ${metric}^{\ast}$ (e.g., mIoU* and PQ*) that reports the average of the perspective metric and fisheye metric. To verify that PanopticNeRF-360 is able to improve the underlying geometry, we further evaluate on perspective views the rendered depth compared to sparse depth maps obtained from LiDAR using Root Mean Squared Error (RMSE) and the ratio of accurate predictions ($\delta_{1.25}$[3, 18].
4.1 Label Transfer on KITTI-360 Dataset
As most baselines are not designed for panoptic label transfer, we first compare the semantic predictions, and then compare the panoptic predictions to 3D-2D CRF.
Semantic Label Transfer: As shown in Table I, Fig. 8 and Fig. 9, our method achieves the highest mIoU and Acc quantitatively and qualitatively. Specifically, compared to 3D-2D CRF, we obtain an absolute mIoU improvement of $2.4\%$ on perspective views and $11.9\%$ on fisheye views. The performance gap widens on the fisheye view as the pseudo-GT is less accurate on fisheye imagery. This observation also highlights the value of our method which is able to provide RGB images and semantic labels at novel viewpoints to enhance the generalization ability of 2D perception models. Despite PSPNet* being finetuned on KITTI-360 which reduces the performance gap, our method outperforms PSPNet* by a large margin. Supervised by the extremely sparse manually annotated GT, Semantic NeRF struggles to produce reliable performance. Using pseudo labels of PSPNet* and Tao et al., JacobiNeRF and Panoptic Lifting are capable of denoising and thus improving performance $(67.2\%\rightarrow 68.5\%/69.6\%,46.3\%\rightarrow 52.1\%/52.5\%)$. However, both JacobiNeRF and Panoptic Lifting are inferior in urban scenarios when the input views are sparse. In terms of perspective MC, ours is comparable to 3D-2D CRF and significantly surpasses 2D-to-2D label transfer methods. While our method slightly lags behind 3D Point + GC in terms of MC, this can be explained as label consistency is evaluated on sparsely projected 3D points which GC takes as input to generate a dense label map.
Panoptic Label Transfer: We compare against Panoptic Lifting and 3D-2D CRF for panoptic label transfer, as the other baselines do not support distinguishing instances. As Panoptic Lifting is limited to segment instance classes known to pre-trained 2D panoptic segmentation models, it is not capable of segmenting the “building” class. Therefore, we consider two settings for quantitatively measuring the panoptic segmentation metrics: “building” as stuff when comparing with Panoptic Lifting, and “building” as thing when comparing with 3D-2D CRF. Table II shows that PanopticNeRF-360 outperforms the baselines in both cases. As shown in Fig. 10, Panoptic Lifting can not reconstruct instances well in far regions where Mask2Former and their instance assignment strategy perform poorly, yielding a larger gap in terms of the panoptic segmentation metrics. In contrast, PanopticNeRF-360 avoids this issue by leveraging weak 3D labels, consistently producing good performance in both near and far regions. Our proposed method also outperforms the 3D-2D CRF in both things and stuff classes, despite that 3D-2D CRF leverages the same weak 3D labels as input. Visual comparisons are shown in Fig. 10 and Fig. 11. As can be seen, our method gracefully handles over- and under-exposure at buildings, which is a challenge for the 3D-2D CRF, as it projects LiDAR points to reconstruct the intermediate meshes whose quality suffers in these scenarios (see Supplementary).
4.2 Label Transfer on Waymo Dataset
We additionally conduct experiments on the Waymo dataset [50] to showcase PanopticNeRF-360’s generalization ability. We choose two scenes, each consisting of 198 consecutive frames with the front-view camera. During data preparation, we use the KITTI-360 annotation toolkit to label 3D bounding primitives and generate 2D pseudo semantic labels with Mask2Former [13]. This annotation takes around 2 hours for each scene, i.e., the average per-frame labeling time takes only 0.6 minutes. We illustrate the entire set of labeled bounding boxes and camera trajectory in Fig. 13. Ours not only renders high-fidelity appearance (28.28 dB on test views), but also generates high-quality semantic and instance masks on both pre-recorded and novel viewpoints. This demonstrates the applicability of our method.
Method Test View F.T. Data Road Sdwlk Terr Bldg Vegt Car Wall Fence Sky mIoU Acc
Mask2Former Persp. - 93.1 61.3 37.0 88.7 87.7 92.3 31.8 61.4 94.8 72.0 89.3
Mask2Former-ft (coarse) (a) 94.2 66.2 37.5 90.1 89.6 92.8 35.7 63.1 94.3 73.7 90.2
Mask2Former-ft (coarse+ours) (a)+(b) 94.5 71.3 37.9 90.3 89.5 93.3 38.2 64.7 94.5 74.9 90.7
Mask2Former-ft (ours) (b) 94.6 72.1 38.4 90.9 89.3 93.4 40.9 66.8 94.6 75.7 91.1
Mask2Former Fisheye - 84.5 60.7 20.5 80.9 77.1 60.9 26.5 42.3 98.4 61.3 89.1
Mask2Former-ft (coarse) (a) 87.9 64.7 25.3 83.7 89.1 62.3 33.9 43.7 98.1 65.4 89.8
Mask2Former-ft (coarse+ours) (a)+(b) 88.3 68.4 26.5 83.5 88.9 61.9 40.3 44.5 97.7 66.7 90.4
Mask2Former-ft (ours) (b) 88.4 70.3 27.2 84.9 88.4 62.3 43.1 45.8 98.3 67.6 90.8
TABLE III: Quantitative Comparison of Semantic Label on Test Scenes of KITTI-360. (a): 512 coarse pseudo 2D labels (PSPNet*); (b): 1,536 synthesized labels using our method.
@@ 4.2 Label Transfer on Waymo Dataset
TABLE IV: Neural Scene Representation Study over 4 scenes. We bold and underline the two best performing methods, respectively.
4.3 Neural Scene Representation
@@
We further compare iNGP [54], Tri-planes [9] (TensoRF-VM), and MLP [52] as scene representation. Towards this goal, we analyze the results over 4 scenes where ambient conditions vary between each other. For a fair comparison, we run each scene for 30,000 iterations. Please refer to the supplementary for implementation details. Table IV indicates that while the pure MLP demonstrates superior label quality, it falls short in appearance reconstruction. On the other hand, iNGP excels in novel view synthesis ($\sim$3.5dB higher than MLP) and convergence, albeit at the cost of subpar semantic label quality in comparison to the MLP. Our proposed scene feature, a hybrid of MLP and hash grids, achieves a tradeoff between label quality and appearance. Tri-planes fail to represent high-fidelity scene structure both in terms of semantic labels and appearance, especially at boundaries between foreground and background (sky) and in over-exposed regions as grid-based methods are not robust to filter noise and tend to suffer from local minimum. While iNGP is also grid-based, it conditions features on multi-scale levels, thereby potentially aggregating global information, and employs a hash function for non-periodic signal embedding. As depicted in Fig. 15, iNGP contains more noise than ours, and Tri-planes fail to reconstruct the overall label, especially at buildings over-exposed by sunlight.
@@ 4.4 Synthesized Labels for Perception Models
To validate that the synthesized novel view labels are important for autonomous driving and robotics applications, we fine-tune Mask2Former [13] pre-trained on CityScape [15] using our synthesized semantic labels on the KITTI-360. We fine-tune it using our synthesized semantic labels on the KITTI-360. Specifically, we divide 10 test scenes into 8 for training and 2 for test. For each scene, we render 3 semantic maps at each of 64 trajectory points using the left perspective camera: one in the original direction, one rotated horizontally 30° to the left, and another rotated horizontally 30° to the right, resulting in a total of 1,536 training samples. We further compare Mask2Former-ft (ours) with two additional variants: one fine-tuned solely on coarse semantic labels that we use as pseudo 2D GTs (coarse), and another trained on a combination of coarse and synthesized images (coarse + ours). During fine-tuning, we format the data to match CityScape’s and exclude categories not present in CityScape. We fine-tune all the models for 5 epochs with a batch size of 8, using a learning rate of 0.0001 and AdamW optimizer. As illustrated in Table III and Fig. 14, Mask2Former-ft (ours) significantly outperforms the original (+3.7 mIoU on perspective views, and +6.3 mIoU on fisheye views), especially on Sidewalk and Wall. It is also reasonable that Mask2Former-ft (ours) outperforms the other two finetuning baselines as our synthesized labels significantly surpass the base 2D pseudo labels. This improvement demonstrates the effectiveness of our synthesized labels in enhancing the perception model’s performance. Moreover, the larger improvement on the fisheye views further verifies that existing perception models degenerates on unseen viewpoints and demonstrates the importance of being able to synthesizing labels with large viewpoint changes.
@@ 4.5 Ablation Study
5 Conclusion
We present PanopticNeRF-360 that infers in 3D space and renders omnidirectional per-pixel semantic and instance labels for 3D-to-2D label transfer. By unifying coarse 3D bounding primitives and noisy 2D semantic predictions, PanopticNeRF-360 is capable of performing mutual enhancement of geometry and semantics compared to naïve joint optimization of geometry and semantics. Specifically, it improves the underlying scene geometry given sparse input views leveraging label-guided geometry optimization, while concurrently resolving label noise based on improved geometry. Moreover, it enables label synthesis at a large range of novel viewpoints, including panoramic perspectives. We posit that our method marks a significant step towards improving data annotation efficiency while delivering a consistent, continuous 3D panoptic representation.